# Formal Definition and Implementation of Reproducibility Tenets for Computational Workflows

### v1

@@ 5.3 Demonstrative Lowpass Workflow
Figure 2: Generic lowpass filter workflow. The filter component is swapped out for specific implementations, but the rest of the workflow remains the same. Boxes with blue bars represent tasks, the RAM logos represent data storage in memory. The final result is stored in a file.
The workflow used to demonstrate the implemented functionality generates a pure sine wave, injects random Gaussian noise, and then attempts to filter the result with a lowpass filter (based on convolution with a Hann window). While structurally simple, each component is highly configurable and moreover, we implement the core filter component with four different methods; a NumPy [47] pointwise brute-force convolution filter for reference, a NumPy FFT filter, a FFTW [48] FFT filter and PyCuda [49] GPU based FFT filter. Figure 2 contains a single rendering of this workflow. This workflow highlights numerical differences between FFT implementations, a fundamental operation in astronomical sciences and wider scientific computing.
Figure 3: Workflows including the filter task and NCC calculation. Symbols are as in Figure 2. The hard disk icons represent files.
In addition to the basic workflow, we construct an extended workflow that computes the Normalized Cross-Correlation NCC [50] between the original noisy signal and the recovered filtered signal divided by the length of the signal. The resulting metric effectively gives the probability that any filtered sample matches its corresponding sample in the original signal. Figure 3 contains a single rendering of this workflow using the DALiuGE graphical workflow editor. This metric is relative, and therefore, we use the length of the signal to determine the numerical precision required to compare our methods as:
@@
Our signal length is $512$ in all trials, so we require three digits of precision in our NCC metric. We execute each of the basic and extended workflows 10 times, changing the random seed for the noise generator to the trial number in both cases.
We now outline our example use cases for each tenet:
@@
We compare the rerun hashes of each standard workflow. Successful reruns mainly establish trust in the workflow management system itself. While each workflow uses different components, they are all implemented in Python and therefore should be reruns.
@@
We compare the repeat hashes of each extended workflow trial. Trials that match can effectively have their results bundled together in concert in making a subsequent claim, such as the efficacy of a filter in this case.
@@
We compare the recompute hashes of each workflow trial. Without an identical computing environment, one would not expect any workflow to be a recomputation. Establishing that workflows are not recomputations guarantees differences between them.
@@
For each repeat trial, we compare the data hash of the signal and window data components; the signal components should differ, but the window should remain identical. By comparing the reproduce signature for each trial, we see if each type of filer consistently filters to the same quality and whether this quality differs between each implementation. This test exhibits how building result post-processing into workflows directly allows the precise testing of highly abstract ideas (e.g., the performance of a filter implementation).
@@
For each type of replication, we compare the hashes of the extended workflow trials. One would expect some repeat trials of the same method to be scientific replicas since the performance of a particular filter should be consistent. A computational replica would be surprising, but establishing total replication on different computing hardware verifies our claims.
6 Demonstration
To demonstrate our methodology as a practical means of comparing workflow designs and implementations, we provide example workflow executions designed to highlight our tenets and real-world examples of where they would apply to astronomical interferometric data processing. All tenet signatures for all referenced workflow trials are available online [51]. Running multiple trials of each workflow, changing only the random noise introduced to the base signal, and comparing our tenet hashes provides a straightforward way to explore what they uncover. Table 4 in appendix A contains the truncated hashes of all plain workflow executions, omitted here for brevity. Table 5 contains the truncated hashes of all NCC test workflow executions, also omitted here for brevity.
For our lowpass filter workflow, we find that although each workflow uses a different library implementation, all trials show the same rerun signature. Repeat trials, differing only in their component parameterization, are identified as such; repeat hashes are the same between implementation methods but differ from their equivalent rerun signatures. In the complete cross-correlation trials, most, but not all, repeat executions are reproductions, indicating that regardless of input data specifics, the filtering component recovers an equivalent amount of signal. The slight variation in signatures indicates that some methods are less effective for some inputs. Upon further inspection, the NCC measure differs by 0.001, which is replicated across all methods on the same input trials. By singling out the input signal data components and comparing their reproduction data hashes, we can confirm that each repeated trial operates on different data. Moreover, despite differing implementations, many cross-correlation repeat trials are reproductions between implementations, demonstrating that each method performs equally well as a filter despite differences in the specific data results. No trials are recomputations, which is to be expected since the runtime trace of each workflow will vary slightly even on the same machine. Scientific replicas single out trials with identical filtering performance and method, formally asserting the equivalence between the FFTW, Cuda FFT, Numpy FFT, and Numpy Pointwise workflows. No trials are computational replicas. Trials matching in workflow and reproductions are total replicas.
@@ 7 Conclusion
Computational workflows power contemporary data-intensive sciences. Workflows play an increasingly important role in astronomical sciences in particular. To build scientific trust in these complex software systems, research engineers adopt industry-standard software development methodologies to implement and test the software. We provide an additional methodology for testing and verifying consistency between and across execution runs using either the same or varying implementations of specific algorithms. We construct a hash-graph-based signature mechanism termed BlockDAGs to characterize workflow executions with respect to these standards and add the resulting hash keys to the provenance information of individual runs. Our demonstrative lowpass filter example shows that this method asserts the scientific equivalence of four different implementations of the same basic workflow while simultaneously identifying the output-data differences between them. Workflow management systems guide workflow execution and design, and our method to test the scientific equality between workflow executions allows us to verify that these systems deliver on their promised benefits. Additional considerations, such as adding formal provenance data to the component hashes and assessing the impact of hardware failure on large scientific workflow instances, would further verify the efficacy of our method. Implementing our BlockDAG mechanism in other workflow management systems would allow for workflow verification between workflow management systems. By providing a test-driven method to determine the scientific equivalence between workflow executions we provide verifiable evidence that next-generation science instruments, such as the SKA produces high quality, reproducible results.
@@ Appendix A Workflow trial signatures
FFT-np-0 8619d 6bc99 76aac 8fd08 faf14 eac3f 07d3a
FFT-np-1 8619d 6bc99 326f1 bf29b 52ae2 3748e 9b208
FFT-np-2 8619d 6bc99 c493b bf29b 52ae2 64fa3 9b208
FFT-np-3 8619d 6bc99 09956 8fd08 faf14 dc375 07d3a
FFT-np-4 8619d 6bc99 e15ba bf29b 52ae2 2a5df 9b208
FFT-np-5 8619d 6bc99 c261c bf29b 52ae2 ef2c5 9b208
FFT-np-6 8619d 6bc99 a3cc5 bf29b 52ae2 a59c6 9b208
FFT-np-7 8619d 6bc99 7295c bf29b 52ae2 72836 9b208
FFT-np-8 8619d 6bc99 a03e0 8fd08 faf14 3131e 07d3a
FFT-np-9 8619d 6bc99 88ff5 bf29b 52ae2 6d378 9b208
FFT-cuda-0 8619d 85f98 09b82 8fd08 faf14 9343a 7ecb0
FFT-cuda-1 8619d 85f98 0ebd4 bf29b 52ae2 2df4d 6f8a9
FFT-cuda-2 8619d 85f98 641dc bf29b 52ae2 78071 6f8a9
FFT-cuda-3 8619d 85f98 7e4bf 8fd08 faf14 d7105 7ecb0
FFT-cuda-4 8619d 85f98 59a07 bf29b 52ae2 4f18d 6f8a9
FFT-cuda-5 8619d 85f98 fb74a bf29b 52ae2 46996 6f8a9
FFT-cuda-6 8619d 85f98 6b8ea bf29b 52ae2 25474 6f8a9
FFT-cuda-7 8619d 85f98 8581f bf29b 52ae2 a9abf 6f8a9
FFT-cuda-8 8619d 85f98 900e1 bf29b 52ae2 93555 6f8a9
FFT-cuda-9 8619d 85f98 024b7 bf29b 52ae2 1dfe3 6f8a9
@@
Pointwise-np-ncc-0 3cabf deb85 d5556 4fa43 f9615 f321d ca359
Pointwise-np-ncc-1 3cabf deb85 117c4 9ae2d bfc4b 1e3cf 4d937
Pointwise-np-ncc-2 3cabf deb85 03d3c 26e2a 8b785 17ca2 7fb60
Pointwise-np-ncc-3 3cabf deb85 3bc5e 0d4a1 6648d d0034 687bf
Pointwise-np-ncc-4 3cabf deb85 58e17 13c75 33366 c433b 6d1fe
Pointwise-np-ncc-5 3cabf deb85 e1a4b 2d0d5 4ce42 56002 b4138
Pointwise-np-ncc-6 3cabf deb85 284d6 b345f 8c086 fa69b 589bb
Pointwise-np-ncc-7 3cabf deb85 a5f11 dd944 5959c 882e0 58b64
Pointwise-np-ncc-8 3cabf deb85 c8e02 7b758 37c0d 3f3c5 07808
Pointwise-np-ncc-9 3cabf deb85 d9d6b 0da65 7c3fb 29eb0 b6bcc
FFT-fftw-ncc-0 3cabf 2f7f9 7cb68 3449a 837f2 70370 7549c
FFT-fftw-ncc-1 3cabf 2f7f9 080d7 928a7 aa7a0 45b3a 5b6a9
FFT-fftw-ncc-2 3cabf 2f7f9 24963 59f82 305bd 36c5f c7ce8
FFT-fftw-ncc-3 3cabf 2f7f9 6fa7a efb06 c2896 ef68f 2e278
FFT-fftw-ncc-4 3cabf 2f7f9 4b180 b3742 c0e64 5b7bb 4eb40
FFT-fftw-ncc-5 3cabf 2f7f9 3e519 f07ec 65a85 0a35f 00308
FFT-fftw-ncc-6 3cabf 2f7f9 0c776 501eb b3fae 8d837 3d6b0
FFT-fftw-ncc-7 3cabf 2f7f9 6d9bc 6ebb8 d4f24 60547 f4bd6
FFT-fftw-ncc-8 3cabf 2f7f9 4019e b9cf6 fb869 2c487 d09bc
FFT-fftw-ncc-9 3cabf 2f7f9 91d11 5ce91 3be4d b2ae1 9c36d

### v2

@@ 6.1 Demonstrative Lowpass Workflow
We first present two workflows implemented in DALiuGE and designed in Eagle [60], a graphical workflow designer for DALiuGE. The first workflow generates a pure sine wave, injects random Gaussian noise, and then attempts to filter the result with a lowpass filter (based on convolution with a Hann window). While structurally simple, each component is highly configurable and we implement the core filter component with four different methods; a NumPy [61] pointwise brute-force convolution filter for reference, a NumPy Fast Fourier Transform (FFT) filter, a FFTW [62] FFT filter and PyCuda [63] GPU based FFT filter. Figure 15 contains a single rendering of this workflow. This workflow highlights numerical differences between FFT implementations, a fundamental operation in astronomical sciences and wider scientific computing.
Figure 15: Generic lowpass filter workflow. The filter component is swapped out for specific implementations, but the rest of the workflow remains the same. Circles with Python logos represent Python tasks, and the RAM logos represent data storage in memory. The file icon represents a result written to storage and is accessible after workflow execution.
Figure 16: Workflows including the filter task and NCC calculation. Symbols are as in Figure 15.
In addition to the basic workflow, we construct a second workflow that computes the Normalized Cross-Correlation NCC [64] between the original noisy signal and the recovered filtered signal divided by the length of the signal. The resulting metric effectively gives the probability that any filtered sample matches its corresponding sample in the original signal. Figure 16 contains a single rendering of this workflow using EAGLE, the DALiuGE graphical workflow editor [60]. This metric is relative, and therefore, we use the length of the signal to determine the numerical precision required to compare the different methods as:
@@
The signal length is $512$ in all trials, so we require three digits of precision in the NCC metric. We execute each basic and extended workflow ten times, changing the random seed for the noise generator to the trial number in both cases.
To demonstrate this methodology as a practical means of comparing workflow designs and implementations, we provide example workflow executions designed to highlight the reproducibility tenets and real-world examples of where they would apply to astronomical interferometric data processing. All tenet signatures for all referenced workflow trials are available online [65]. Running multiple trials of each workflow, changing only the random noise introduced to the base signal, and comparing tenet hashes provide a straightforward way to explore what they uncover. Table 10 in A contains the truncated hashes of all plain workflow executions, omitted here for brevity. Table 11 contains the truncated hashes of all NCC test workflow executions, omitted here for brevity.
We now outline example use cases for each tenet:
@@
We compare the Rerun hashes of each standard workflow execution and of each of the NCC workflows. Successful Reruns intend to establish trust in the workflow management system itself, as mentioned previously. While each workflow uses different components, they are all implemented in Python and should all be Reruns, which we find quite straightforwardly to be the case, shown by the identical Rerun hashes.
@@
We compare the Repeat hashes of each extended workflow trial and find that the Repeat hashes of each filter method match in their signature. Trials that match can effectively have their results bundled together to make a subsequent claim. We can confidently conclude that each NCC workflow execution for each filter method did execute the same workflow components, independent of the data they ingested or output.
@@
We can compare the Recompute hashes of each workflow trial and find that all trials have a different Recomputation hash. Without an identical computing environment, one would not expect any workflow execution here to be a Recomputation. As defined, demonstrating workflow Recomputation is an extraordinarily difficult task.
@@
For each Repeated trial, that is, bundling signatures based on filter implementation, we see that the individual filter workflows all produce different reproduction signatures. This is to be expected since each method receives a different signal as input owing to the random noise injected, producing subtly different outputs. However, when comparing the Repeated trials for the NCC workflows, we see that some, but not all, workflow executions are Reproductions. Since the final output of the NCC workflows is effectively a reconstruction error and despite the input being different for each trial, all executions fall into one of two signatures: ‘8fd08’ or ‘bf29b’. Upon further inspection, we find a slight but potentially impactful difference of 0.001. This test exhibits how building results post-processing into workflows directly allows for the precise testing of highly abstract ideas (e.g., the performance of a filter implementation).
@@
For each type of replication, we compare the hashes of the extended workflow trials. One would expect some Repeat trials of the same method to be scientific replicas since the performance of a particular filter should be consistent. A computational replica would be highly surprising, and establishing total replication verifies that a particular method produces a particular result. We find that workflows that are Reproductions are also scientific and total Replications. By formally demonstrating that different filtering executions produce different results, regardless of what filtering method is used, but also showing that the final error in filtering is consistent across each method tested, we can assert with confidence an equivalence between the FFTW, Cuda FFT, Numpy FFT, and Numpy Pointwise workflows.
@@ 7 Conclusion
Computational workflows power contemporary data-intensive sciences, and play an increasingly important role in astronomical sciences in particular. Research engineers adopt industry-standard software development methodologies to build scientific trust in these complex software systems to implement and test the software. We provide an additional methodology for testing and verifying consistency between and across execution runs using either the same or varying implementations of specific algorithms. We formally define seven reproducibility tenets based on a generic workflow UML model. We then consider specific provenance information required to satisfy each definition and construct a hash-graph-based signature mechanism termed BlockDAGs. We provide a concrete example implementation of this mechanism in DALiuGE connecting the UML scale-agnostic workflow model, reproducibility tenet definitions, which inform provenance information collection, and workflow signature mechanism together. A demonstrative lowpass filter example shows that this method asserts the scientific equivalence of four implementations of the same basic workflow while simultaneously identifying their output-data differences. Workflow management systems guide workflow execution and design, and this method of testing the scientific equality between workflow executions allows us to verify that these systems deliver their promised benefits. Additional considerations, such as adding formal provenance data to the component hashes and assessing the impact of hardware failure on large scientific workflow instances, would further verify the efficacy of this method. Implementing a BlockDAG mechanism in other workflow management systems would allow for workflow verification between workflow management systems. By providing a test-driven method to determine the scientific equivalence between workflow executions, we provide verifiable evidence that next-generation science instruments, such as the SKA, produce high-quality, reproducible results.
@@ Appendix A Workflow trial signatures
FFT-np-0 3cabf dfbfa 6d111 bf3ac c7cca 3954f 7f973
FFT-np-1 3cabf dfbfa 9c1bc c6a23 dc639 0562e e9632
FFT-np-2 3cabf dfbfa 287a0 00598 8beb1 48677 6db23
FFT-np-3 3cabf dfbfa dc330 c52dc 39eb8 62956 d9c33
FFT-np-4 3cabf dfbfa a33aa 19172 819ce 90cb3 4d5a6
FFT-np-5 3cabf dfbfa 3d34d 3880c d23fe 96fc3 337af
FFT-np-6 3cabf dfbfa 57c3b 327c7 a64b6 a6687 3894f
FFT-np-7 3cabf dfbfa f2317 d196f 50607 ba3ec b47f6
FFT-np-8 3cabf dfbfa 1732f 2e491 8d5ab 856a5 eff00
FFT-np-9 3cabf dfbfa 8ed58 5c0bf abff8 8ed1b bcc57
FFT-cuda-0 3cabf 1df3e 4e149 77f90 4689c 9bfa4 6eb83
FFT-cuda-1 3cabf 1df3e 0fcc0 f1f8c 200d5 16eaa b06e0
FFT-cuda-2 3cabf 1df3e bddd5 f9381 99da5 b0cb3 23392
FFT-cuda-3 3cabf 1df3e 8aae0 f5a85 7320e 6bc92 5deaa
FFT-cuda-4 3cabf 1df3e ac2fb 0124a 5d9b9 2ebe6 0f4c9
FFT-cuda-5 3cabf 1df3e 6e775 55d9b 261d5 375ca f57c3
FFT-cuda-6 3cabf 1df3e 9994d d1cc2 6c0c1 d5dcd c497e
FFT-cuda-7 3cabf 1df3e 4c48b 52427 fd3f6 4ba05 e4b81
FFT-cuda-8 3cabf 1df3e 41009 150ef fbbea d777f d71f9
FFT-cuda-9 3cabf 1df3e eb6d0 9f45e f7215 ca7f8 55be9
@@
Pointwise-np-ncc-0 8619d c7529 78d51 8fd08 faf14 998cf f353b
Pointwise-np-ncc-1 8619d c7529 57d1d bf29b 52ae2 c7526 fcf65
Pointwise-np-ncc-2 8619d c7529 61f9a bf29b 52ae2 95914 fcf65
Pointwise-np-ncc-3 8619d c7529 2d4d2 8fd08 faf14 2094b f353b
Pointwise-np-ncc-4 8619d c7529 8128d bf29b 52ae2 654a3 fcf65
Pointwise-np-ncc-5 8619d c7529 c4615 bf29b 52ae2 52773 fcf65
Pointwise-np-ncc-6 8619d c7529 8772c bf29b 52ae2 5d4fd fcf65
Pointwise-np-ncc-7 8619d c7529 576e6 bf29b 52ae2 faf9e fcf65
Pointwise-np-ncc-8 8619d c7529 2e510 8fd08 faf14 0fc59 f353b
Pointwise-np-ncc-9 8619d c7529 b5418 bf29b 52ae2 4b7b5 fcf65
FFT-fftw-ncc-0 8619d 6bc99 76aac 8fd08 faf14 eac3f 07d3a
FFT-fftw-ncc-1 8619d 6bc99 326f1 bf29b 52ae2 3748e 9b208
FFT-fftw-ncc-2 8619d 6bc99 c493b bf29b 52ae2 64fa3 9b208
FFT-fftw-ncc-3 8619d 6bc99 09956 8fd08 faf14 dc375 07d3a
FFT-fftw-ncc-4 8619d 6bc99 e15ba bf29b 52ae2 2a5df 9b208
FFT-fftw-ncc-5 8619d 6bc99 c261c bf29b 52ae2 ef2c5 9b208
FFT-fftw-ncc-6 8619d 6bc99 a3cc5 bf29b 52ae2 a59c6 9b208
FFT-fftw-ncc-7 8619d 6bc99 7295c bf29b 52ae2 72836 9b208
FFT-fftw-ncc-8 8619d 6bc99 a03e0 8fd08 faf14 3131e 07d3a
FFT-fftw-ncc-9 8619d 6bc99 88ff5 bf29b 52ae2 6d378 9b208